OpenStudy (anonymous):

Your test scores in science are 83 and 87. What possible scores can you earn on your next test to have a test average between 85 and 90, inclusive?

OpenStudy (anonymous):

@confluxepic

OpenStudy (confluxepic):

It depends. @KALIAN

OpenStudy (anonymous):

watch mean ?

OpenStudy (texaschic101):

(83 + 87 + x) / 3 = 85 (170 + x) / 3 = 85 170 + x = 255 x = 255 - 170 x = 85 so he needs an 85 to make an 85 (83 + 87 + x) /3 = 90 (170 + x) / 3 = 90 170 + x = 270 x = 270 - 170 x = 100 so he needs a 100 to make a 90

OpenStudy (texaschic101):

so I guess the answer would be : at least 85 but no more then 100

OpenStudy (confluxepic):

It has to be over 85.

OpenStudy (confluxepic):

About 90.

OpenStudy (anonymous):

so 85 to what ?

OpenStudy (texaschic101):

oh..between 85 and 90....answer would be any number between 85 and 100

OpenStudy (anonymous):

Ok Thanks guys :) Who deserves the medal . . .

OpenStudy (confluxepic):

I gave @texaschic101 one.

OpenStudy (texaschic101):

your choice

OpenStudy (confluxepic):

I doesn't really matter.

OpenStudy (anonymous):

hmmm

OpenStudy (confluxepic):

I don't really care.

OpenStudy (texaschic101):

I will give confluxepic one

OpenStudy (confluxepic):

Oh well.

OpenStudy (confluxepic):

Problem solved.

OpenStudy (texaschic101):

:) happy,happy,happy

OpenStudy (anonymous):

k i give chic one, chic gives flux one

OpenStudy (anonymous):

Rok da Luv !